Can someone please help me with this sum its hard and i have tried to attempt it a few times this question is really tricky! part of the question is in fraction for .this is a simulatenous linear equations
2x+3y=-30
y=2/3x-6
My answer:
2x+3(2/3x-6)=-30

Answer by mananth(16949)

2x+3y=-30
y= (2/3)x-6
In such problems eliminate the denominator by multiplying by the LCD.
LCD here is 3
3y= 2x-18
substitute 3y in the first equation
2x+3y=-30
2x+2x-18 = -30
4x-18 =-30
+18
4x=-30+18
4x= -12
/4
x= -3
Plug the value of x in any equation above to get y
we get y=-8
(-3,-8)